A particle is projected ...at that time its kinetic energy was KE1......and at the top point where vertical velocity is 0 its kinetic energy was KE2 .......


The ratio of KE2 : KE1 is 3:4 ........ find the difference in their velocities ......




 



































Dear ankit


let velocity at the begineeing ii u and angle of projection is θ.


KE1 =1/2 mu2


KE2 =1/2 m(ucosθ)2


given


KE2 : KE1 is 3:4


so cos2θ  =3/4


θ =30


difference in their velocity =u-ucosθ


                                     =u-u√3/2


                                     =u(1-√3/2)


 


sir but answer of this is u/2 ....



Grade:

2 Answers

chaitanya pansare
14 years ago

then might be the ques. is asking the difference in vertical velocities which comes out to be usinθ-0 i.e. u/2-0=u/2

Badiuddin askIITians.ismu Expert
14 years ago

Dear Ankit Shukla

last answer was on the basis of difference in velocity magnitude.

I think you want difference in velocity vector

Initial velocity =ucosΘ i + usinΘ j

final velocity =ucosΘ i

 

difference in velocity =ucosΘ i + usinΘ j -ucosΘ i

                             =usinΘ j

                            =u sin30 j

                            =u/2 j

magnitude            =|u/2 j|

                           =u/2

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin